hartnn (hartnn):

Find the inverse Z-Transform of - f(z)=(z+2)/(z^2-2z+1) ,|z|>1

hartnn (hartnn):

So i have \(f(z)=\dfrac{1}{(z-1)}+\dfrac {3}{(z-1)^2} = \dfrac{z^{(-1)}}{(1-z^{(-1)} )}+\dfrac{3 z^{(-2)}}{(1-z^{(-1)} )^2} \)

hartnn (hartnn):

I know the INV Z Transform of 1st term, but how about the 2nd term ??

hartnn (hartnn):

or is there any other way to do this ?

hartnn (hartnn):

convolution ?

OpenStudy (anonymous):

What if you use Partial Fraction for \(\large \frac{X(z)}{z}\)

OpenStudy (anonymous):

Sorry that is F(z) there..

hartnn (hartnn):

tried that..no use..

OpenStudy (anonymous):

Let me try this. :)

OpenStudy (anonymous):

I don't know this seriously as In Signals, I have not reached here, but I see what I can do in this.. :)

OpenStudy (usukidoll):

what subject is this?

hartnn (hartnn):

Z transform in Maths ...

OpenStudy (usukidoll):

:/

OpenStudy (usukidoll):

do you know pdes?

hartnn (hartnn):

Partial DE's ...yes, but I don't think that should be used here...

OpenStudy (usukidoll):

it's because I'm in that course and I don't understand what this person is talking about. you can read my question here http://math.stackexchange.com/questions/922338/find-a-solution-of-u-xxu-yy-axbyc-for-given-real-constants-a-b-and?noredirect=1#comment1904098_922338

OpenStudy (anonymous):

Let X(z) = F(z)/z

OpenStudy (anonymous):

\[X(z) = \frac{F(z)}{z} = \frac{z+2}{z(z-1)^2} = \frac{2}{z} - \frac{2}{z-1} + \frac{3}{(z-1)^2}\] Check if it is right?

OpenStudy (anonymous):

Did you get the same Hardik when you tried this method?

hartnn (hartnn):

ok, i see where you're going

hartnn (hartnn):

2-2z/(z-1) +3z/(z-1)^2

OpenStudy (anonymous):

Now multiply by z both the sides, you will get F(z) again..

OpenStudy (anonymous):

Yes.. \[Z^{-1}(\frac{z}{(z-1)^2}) = n \cdot u[n]\]

hartnn (hartnn):

\(2\delta [n] -2u[n]+3 n u[n]\) ?

OpenStudy (anonymous):

Yep, according to me..

hartnn (hartnn):

anyways, i'll ignore wolf and go with your method/answer . thanks! :)

OpenStudy (anonymous):

But if n is greater than 0, then \(\delta[n]\) will be or not?

hartnn (hartnn):

its not unilateral Z-Transform, n can be negative too

OpenStudy (anonymous):

Okay.. Does Long Division Method apply here??

OpenStudy (anonymous):

I think, I can help you till here only because my knowledge is very limited here, I am on Convolution now, then Fourier Series, Transform, then Laplace and then Z-transform will come.. :( I want to go in this sequence in which they have their origin, so still a long journey to cover.. :)

hartnn (hartnn):

good luck!

OpenStudy (anonymous):

Thanks.. :)

OpenStudy (anonymous):

Till then clear your concept on this, I will be in comfort then.. :P

hartnn (hartnn):

sure!

Miracrown (miracrown):

 and what we can do is to write the Z inverse transform for 1/(1-z^-1) as you did and then we square it like above ^